OpenStudy (anonymous):

Use the qoutient of powers property to simplify. 4^2/4, 8^3/8^5, 5^3/5^5, c^4/c^6.

OpenStudy (anonymous):

4^2/4=4^(2-1)=4 8^3/8^5=1/8^2 5^3/5^5=1/5^2 1/c^2